OpenStudy (anonymous):

Find the solution to the system algebraically. Leave your answers in fraction form. Write your answers as ordered pairs using parentheses. If there are two solutions, separate the two solutions with a comma. Do not put a space between them. y = -3 y = x^2 - 4

OpenStudy (texaschic101):

y = -3 now sub -3 in for y in the other equation -3 = x^2 - 4 x^2 - 4 + 3 = 0 x^2 - 1 = 0 x^2 = 1 x = sq rt 1 x = 1 check.. y = x^2 - 4 -3 = 1^2 - 4 -3 = 1 - 4 -3 = -3 (correct) x = 1 and y = -3 or (1,-3)

OpenStudy (anonymous):

Thank you